# Numerically “exact” charge transport dynamics in a dissipative electron–phonon model rationalizing the success of the transient localization scenario

### v1

@@ II.6 Model parameters
We focus on the slow-phonon ($\omega_{0}/J=0.044$), high-temperature ($T/J=0.175$ or $T/\omega_{0}\approx 4$), intermediate-interaction ($\lambda=0.336$) regime, which is relevant for anisotropic carrier transport in organic molecular crystals. [23, 3] The damping parameter $\gamma_{0}$ in Eq. (6) can be estimated from combined molecular-dynamics and quantum-chemistry analyses of the vibrational motions that most strongly modulate carrier transport along the maximum-conductivity direction of rubrene [23] and a discotic liquid crystal. [30] Troisi and coworkers [30, 31, 76] concluded that the carrier–phonon interaction limiting carrier transport in quasi-one-dimensional organic semiconductors can be described by the BO or multimode BO SD. The most prominent mode can be either underdamped $\gamma_{0}/\omega_{0}\sim 0.25$ in Fig. 1(b) of Ref. 23] or overdamped ($\gamma_{0}/\omega_{0}=1.7$ in Ref. 30). Assuming undamped vibrations, we have recently examined carrier transport for different adiabaticity ratios $\omega_{0}/J$, temperatures, and interactions. [24, 25] We do not expect that the damping of vibrational motions change the trends observed upon varying these parameters. Therefore, here we mostly concentrate on understanding the effect of damping on transport dynamics, which has not been widely explored so far.
@@ II.7 Quantities describing carrier transport
which determines the instantaneous power-law growth of $\Delta x^{2}(t)$ with time, $\Delta x^{2}(t)\propto t^{\alpha(t)}$. We are mostly interested in how the details of the crossover between short-time ballistic transport, for which $\alpha(t)\approx 2$, and long-time diffusive transport, for which $\alpha(t)$ approaches unity, change upon varying the damping rate $\gamma_{0}$. In experiments, transport dynamics are usually inferred from the dynamical-mobility profile
@@ III Technical details
The parameter regime described in Sec. II.6 is challenging because the DEOM [Eqs. (27) and (20)] have to be solved on long chains, as implied by the small adiabaticity ratio, and truncated at a moderate maximum depth $D$ dictated by the combination of high temperature and intermediate interaction. Up to now, the (bosonic) dissipaton algebra has been used to study hybrid system–bath dynamics only in few-level systems (monomers and dimers). [72, 71, 58, 57, 73] Meanwhile, the applications of the (bosonic) DEOM formalism to systems with somewhat larger number of electronic levels focused on purely electronic dynamics. [77, 78, 79, 80]
To the best of our knowledge, our study is among the first studies exploiting the (bosonic) dissipaton algebra to compute a finite-temperature real-time correlation function of a mixed system–bath operator in a many-level system. Our computations are made possible by formulating the DEOM in momentum space, [62, 24] in which the DDO $\iota_{\mathbf{n}}^{(n)}(t)$ in Eq. (20) has only $N$ (instead of $N^{2}$) nonzero matrix elements $\langle k|\iota_{\mathbf{n}}^{(n)}(t)|k+k_{\mathbf{n}}\rangle$, and similarly for $\langle k|\sigma_{\mathbf{n}}^{(n)}(\tau)|k+k_{\mathbf{n}}\rangle$ in Eq. (27). In Secs. III.1III.3, we discuss further technical details needed to gain confidence in our DEOM results.
@@ III.1 DEOM closing scheme
Figure 2: (a) Real (black) and imaginary (red) parts of the bath correlation function $\mathcal{C}(t)$ computed using $K=2$ (symbols) and $K=15$ (lines) terms in the exponential decomposition in Eq. (11). The time-dependent diffusion constant (b) and the dynamical-mobility profile (c) computed using DEOM with $\Gamma(k,\mathbf{n})=0$ in Eq. (35) (label “TNL”, thin lines) and with the Markovian-adiabatic [Eqs. (36) and (37), label “MA”, thick solid line] and derivative-resum [Eq. (38), label “DR”, dashed line] closing terms with respect to $D$. The inset of panel (b) compares the diffusion-exponent dynamics computed using the same closing schemes. The model parameters are $J=1,\omega_{0}=0.044,\lambda=0.336,T=0.175,$ and $\gamma_{0}=1.7\omega_{0}$, while the maximum DEOM depth is $D=4$.
Ideally, the r-DEOM in Eq. (20) are to be solved using $K\to\infty$ terms in the exponential decomposition in Eq. (11), and considering DDOs up to the maximum depth $D\to\infty$. In practice, both $K$ and $D$ are finite, and the convergence with respect to them has to be checked. The convergence can be enhanced by devising appropriate DEOM closing schemes with respect to $K$ and DD[81, 82, 57]
As we focus on high temperatures, at which $\beta\omega_{0}\ll 1$ and $\beta\gamma_{0}\ll 1$, see Sec. II.6, it is justified to set $K=2$, i.e., to retain only the terms originating from the poles of $\mathcal{J}_{\mathrm{BO}}(\omega)$ in Eq. (11), see Fig. 2(a) and Appendix A. We then employ the Ishizaki–Tanimura closing with respect to KK[83, 84] augmenting the RHS of Eq. (20) by
@@
to the equations at depth $D$. In contrast to existing approaches, [81, 82, 57] which were gauged on few-level systems, the closing term in Eq. (35) does not introduce couplings among either DDOs at depth $D$ or different matrix elements of a DDO at depth $D$. This feature of Eq. (35) guarantees that the computational cost of the closing remains much smaller than that of one step in DEOM propagation. The complex quantity $\Gamma(k,\mathbf{n})$ depends on the details of the closing, i.e., on the approximations employed to solve the equations at depth $D+1$. Neglecting the couplings to DDOs at depth $D+2$, and using the Markovian and adiabatic approximations yields
@@
Following the practice common in models without explicit dissipation, in Eq. (35) we retain only real parts of $\Gamma(k,\mathbf{n})$, i.e., we neglect the closing-induced renormalization of free-oscillation frequencies of DDOs at the terminal layer. We have checked that the imaginary parts of $\Gamma(k,\mathbf{n})$ in Eq. (35) very weakly affect the overall transport dynamics, and in particular the long-time transport coefficient $\mu_{\mathrm{dc}}$. In undamped models, [62, 24] $\mathrm{Re}\>\Gamma_{\mathrm{MA}}(k,\mathbf{n})$ and $\mathrm{Re}\>\Gamma_{\mathrm{DR}}(k,\mathbf{n})$ are meaningful only in the infinite-chain limit $N\to\infty$, in which these can be evaluated analytically. For $\gamma_{0}\neq 0$, $\mathrm{Re}\>\Gamma_{\mathrm{MA}}(k,\mathbf{n})$ and $\mathrm{Re}\>\Gamma_{\mathrm{DR}}(k,\mathbf{n})$ can be computed for any finite $N$, on which they strongly depend. Although analytical expressions for $\mathrm{Re}\>\Gamma_{\mathrm{MA}}(k,\mathbf{n})$ and $\mathrm{Re}\>\Gamma_{\mathrm{DR}}(k,\mathbf{n})$ in the $N\to\infty$ limit are not available for $\gamma_{0}\neq 0$, we find that their long-chain limits are reached for $N$ much larger than typical chain lengths $N\sim 30$ we consider here. In practice, we find that computing Eqs. (37) and (38) on a chain containing $N_{\infty}=(10^{3}-10^{4})N$ sites provides us with the desired long-chain limits (although $N_{\infty}/N$ generally decreases with increasing $\gamma_{0}$).
@@
In Ref. 25, we have checked that both the Markovian-adiabatic [Eqs. (36) and (37)] and derivative-resum [Eq. (38)] schemes yield virtually identical transport dynamics for $\gamma_{0}=0$. Figures 2(b) and 2(c) confirm that the same holds in the overdamped regime $\gamma_{0}/\omega_{0}=1.7$. These also reveal that, for sufficiently strong damping, the so-called time-nonlocal closing, [87] which sets $\Gamma(k,\mathbf{n})=0$ in Eq. (35), also yields stable long-time diffusive dynamics from which low-frequency dynamical mobility can be meaningfully extracted. This stands in sharp contrast to the undamped case, in which computations of long-time transport coefficients crucially require a closing term [25] because these cannot be accessed by extrapolating the results for small nonzero dampings to the zero-damping limit. [68] Although Figs. 2(b) and 2(c) suggest that the closing terms can be omitted if $\gamma_{0}$ is sufficiently large, all our results rely on the propagation of the closed DEOM. Using one and the same closing scheme throughout the physically reasonable range for $\gamma_{0}$ down to $\gamma_{0}=0$ enables a consistent and fair assessment of trends in transport dynamics with varying $\gamma_{0}$.
@@
Furthermore, Figs. 2(b) and 2(c) suggest that our DEOM closing enhances not only the convergence with respect to the maximum depth, but also with respect to the chain length. Namely, the DEOM dynamics with $N=31$ and Markovian-adiabatic or derivative-resum closing are very similar to those employing a $\approx 50\%$-longer chain ($N=45$) and time-nonlocal closing. In the overdamped regime considered here, the diffusion-constant growth on intermediate timescales and the corresponding dynamical-mobility feature at $\omega\approx\omega_{0}$, see Figs. 2(b) and 2(c) for $N=31$, reflect finite-size effects. Meanwhile, analogous features for weak damping are not artifacts due to insufficient chain length, [25] as discussed in greater detail in Sec. III.3.
@@
Finally, we prefer the Markovian-adiabatic [Eqs. (36) and (37)] to the derivative-resum [Eq. (38)] scheme because of its more favorable computational requirements. Namely, the Markovian-adiabatic scheme needs only $N$ quantities $\gamma_{k}$ computed on an $N_{\infty}$-site chain, while the derivative-resum scheme needs a much larger number of quantities $\Gamma_{\mathrm{DR}}(k,\mathbf{n})$, $N$ quantities for each DDO at depth $D$. Although all these quantities can be computed only once, before the r-DEOM propagation, our derivative-resum scheme requires much more computational time and computer memory to compute and store the closing terms than the Markovian-adiabatic scheme. It is for this reason that we employ the Markovian-adiabatic scheme in the following.
@@ III.2 Thermodynamic quantities
Also, the carrier–bath interaction energy
@@
has a nonzero imaginary part, in contrast to the undamped case. [24, 25] Although Fig. 3(a) shows that both $\mathrm{Re}\>\langle H_{S}\rangle$ and $\mathrm{Re}\>\langle H_{S-B}\rangle$ grow quadratically with $\gamma_{0}$ (for sufficiently large $\gamma_{0}$), they are virtually equal to the corresponding values in the undamped case for physically plausible values of $\gamma_{0}$. Figure 3(b) shows that $\mathrm{Im}\>\langle H_{S}\rangle$ and $\mathrm{Im}\>\langle H_{S-B}\rangle$ grow linearly with $\gamma_{0}$, but remain quite small with respect to the corresponding real parts throughout the physically relevant range $\gamma_{0}/\omega_{0}\lesssim 2$. Therefore, the real parts of Eqs. (40) and (41) obtained using the i-DEOM remain representative of the kinetic and interaction energies, respectively.
@@ III.3 Finite-depth and finite-chain effects for weak damping
Figure 4: Dynamics of (a) the diffusion constant $\mathcal{D}$ and (b) the diffusion exponent $\alpha$ for different values of the chain length NN and maximum hierarchy depth $D$. (c) The dynamical-mobility profile $\mathrm{Re}\>\mu(\omega)$ for different values of $N$ and $D$. The insets in (a) and (b) zoom in the dynamics of $\mathcal{D}$ and $\alpha$, respectively, on intermediate-to-long timescales. The model parameters are $J=1,\omega_{0}=0.044,\lambda=0.336,T=0.175$, and $\gamma_{0}/\omega_{0}=0.05$. Note the logarithmic scale on horizontal axes.
@@
Fixing $N=21$ or 31, we observe that transport properties computed for $D=4,5,6$ or $D=4,5$ are qualitatively similar, though there are some quantitative differences. Namely, the growth of $\mathcal{D}(t)$ beyond $\mathcal{D}(t_{\mathrm{min}})$ [Fig. 4(a)], where $t_{\mathrm{min}}\approx 1/\omega_{0}$, the overshoot of $\alpha(t)$ above unity on intermediate timescales [Fig. 4(b)], and the dynamical-mobility enhancement below $\omega_{0}$ [Fig. 4(c)] are more pronounced with increasing $D$. Meanwhile, increasing $D$ does not change either the position or shape of the displaced Drude peak [Fig. 4(c)], while quantitative changes in transport properties upon increasing $D$ are of the order of $10\%$[25] As DEOM computations for $D=5$ (let alone $D=6$) require significantly more computational resources than those employing $D=4$, we opt for $D=4$ as a good compromise between computational effort and accuracy. The relative accuracy with which the OSR [Eq. (42)] is satisfied is $\sim 10^{-4}$ for all depths considered.
@@
Fixing $D=4$ ($D=5$), we find that $\mathcal{D}(t),\alpha(t),$ and $\mathrm{Re}\>\mu(\omega)$ for $N=21$, 31, and 45 ($N=21$ and 31) are qualitatively and to a large extent quantitatively similar. Finite-chain effects on the dynamics of $\mathcal{D}$ and $\alpha$ come into play for $t\gtrsim\pi/\omega_{0}$, see the insets of Figs. 4(a) and 4(b). We observe that increasing $N$ renders the increase in $\mathcal{D}(t)$ for $t\gtrsim t_{\mathrm{min}}$, the overshoot of $\alpha(t)$ above unity, and the dynamical-mobility enhancement below $\omega_{0}$ somewhat less pronounced. One might thus object to the relevance of these features in the long-chain limit. To refute such an objection, we note that the very same features have been observed in fully quantum [90] or quantum–classical [91, 21, 92, 93] approaches to carrier transport in models with local [90, 91] or nonlocal [21, 92, 93] interaction of a carrier with undamped [90, 91, 21, 92, 93] or moderately damped [21] vibrations. In particular, the quantum–classical study in Ref. 21 considered the same adiabaticity $\omega_{0}/J$ ratio as here, and observed the same long-time (low-frequency) features in $\mathcal{D}(t)$ $\mathrm{Re}\>\mu(\omega)$] on chains with hundreds of sites. By assuming classical vibrational dynamics, which is appropriate for small $\omega_{0}/J$ and at the high temperature considered here, quantum–classical methods manage to reach what would be the $D\to\infty$ limit in the DEOM language. Keeping in mind that increasing $D$ beyond 4 renders the long-time (low-frequency) features more pronounced, their existence for small $\gamma_{0}$ can be considered as firmly established.
@@
The above observations strengthen the conclusions we reached in the disipationless case, [25] firmly establishing the transient nature of the super-to-subdiffusive crossover and the low-frequency enhancement in the carrier’s optical response in the case of weakly damped vibrations moderately coupled to the carrier.

### v2

@@ II.6 Model parameters
We focus on the slow-phonon ($\omega_{0}/J=0.044$), high-temperature ($T/J=0.175$ or $T/\omega_{0}\approx 4$), intermediate-interaction ($\lambda=0.336$) regime, which is relevant for anisotropic carrier transport in organic molecular crystals. [23, 3] The damping parameter $\gamma_{0}$ in Eq. (6) can be estimated from combined molecular-dynamics and quantum-chemistry analyses of the vibrational motions that most strongly modulate carrier transport along the maximum-conductivity direction of rubrene [23] and a discotic liquid crystal. [30] The net effect of these motions can be conveniently approximated using few BO modes, the most prominent of which can be either underdamped $\gamma_{0}/\omega_{0}\sim 0.25$ in Fig. 1(b) of Ref. 23] or overdamped ($\gamma_{0}/\omega_{0}=1.7$ in Ref. 30). Assuming undamped vibrations, we have recently examined carrier transport for different adiabaticity ratios $\omega_{0}/J$, temperatures, and interactions. [24, 25] We do not expect that the damping of vibrational motions will change the trends observed upon varying these parameters. Therefore, here we mostly concentrate on understanding the effect of damping on transport dynamics, which has not been widely explored so far.
@@ II.7 Quantities describing carrier transport
which determines the instantaneous power-law growth of $\Delta x^{2}(t)$ with time, $\Delta x^{2}(t)\propto t^{\alpha(t)}$. We are mostly interested in how the details of the crossover between short-time ballistic transport, for which $\alpha(t)\approx 2$, and long-time diffusive transport, for which $\alpha(t)$ approaches unity, change upon varying the damping rate $\gamma_{0}$. This ballistic-to-diffusive crossover exhibits periods of superdiffusive $\alpha(t)>1$] and subdiffusive $\alpha(t)<1$] carrier dynamics, and we pay special attention to whether its long-time completion is reached from the super- or subdiffusive side. In experiments, crossover dynamics are usually inferred from the dynamical-mobility profile
@@ III Technical details
The parameter regime described in Sec. II.6 is challenging because the DEOM [Eqs. (27) and (20)] have to be solved on long chains, as implied by the small adiabaticity ratio, and truncated at a moderate maximum depth $D$ dictated by the combination of high temperature and intermediate interaction. Up to now, the (bosonic) dissipaton algebra has been used to study hybrid system–bath dynamics only in few-level systems (monomers and dimers). [73, 55, 60, 59, 74] Meanwhile, the applications of the (bosonic) DEOM formalism to systems with a somewhat larger number of electronic levels focused on purely electronic dynamics. [76, 77, 78, 79]
To the best of our knowledge, our study is among the first studies exploiting the (bosonic) dissipaton algebra to compute a finite-temperature real-time correlation function of a mixed system–bath operator in a many-level system. Our computations are made possible by formulating the DEOM in momentum space, [64, 24] in which the DDO $\iota_{\mathbf{n}}^{(n)}(t)$ in Eq. (20) has only $N$ (instead of $N^{2}$) nonzero matrix elements $\langle k|\iota_{\mathbf{n}}^{(n)}(t)|k+k_{\mathbf{n}}\rangle$ and similarly for $\langle k|\sigma_{\mathbf{n}}^{(n)}(\tau)|k+k_{\mathbf{n}}\rangle$ in Eq. (27). In Secs. III.1III.3, we discuss further technical details needed to gain confidence in our DEOM results.
@@ III.1 DEOM closing scheme
Figure 2: (a) Real (black) and imaginary (red) parts of the bath correlation function $\mathcal{C}(t)$ computed using $K=2$ (symbols) and $K=15$ (lines) terms in the exponential decomposition in Eq. (11). The time-dependent diffusion constant (b) and the dynamical-mobility profile (c) computed using DEOM with $\Gamma(k,\mathbf{n})=0$ in Eq. (35) (label “TNL,” thin lines) and with the Markovian-adiabatic [Eqs. (36) and (37), label “MA,” thick solid line] and derivative-resum [Eq. (38), label “DR,” dashed line] closing terms with respect to $D$. The inset of panel (b) compares the diffusion-exponent dynamics computed using the same closing schemes. The model parameters are $J=1,\omega_{0}=0.044,\lambda=0.336,T=0.175,$ and $\gamma_{0}=1.7\omega_{0}$, while the maximum DEOM depth is $D=4$.
Ideally, the r-DEOM in Eq. (20) is to be solved using $K\to\infty$ terms in the exponential decomposition in Eq. (11), and considering DDOs up to the maximum depth $D\to\infty$. In practice, both $K$ and $D$ are finite, and the convergence with respect to them has to be checked. The convergence can be enhanced by devising appropriate DEOM closing schemes with respect to $K$ and DD[80, 81, 59]
As we focus on high temperatures, at which $\beta\omega_{0}\ll 1$ and $\beta\gamma_{0}\ll 1$ (see Sec. II.6), it is justified to set $K=2$, i.e., to retain only the terms originating from the poles of $\mathcal{J}_{\mathrm{BO}}(\omega)$ in Eq. (11) [see Fig. 2(a) and Appendix A]. We then employ the Ishizaki–Tanimura closing with respect to KK[82, 83] augmenting the RHS of Eq. (20) by
@@
to the equations at depth $D$. In contrast to existing approaches, [80, 81, 59] which were gauged on few-level systems, the closing term in Eq. (35) does not introduce couplings among either DDOs at depth $D$ or different matrix elements of a DDO at depth $D$. This feature of Eq. (35) guarantees that the computational cost of the closing remains much smaller than that of one step in DEOM propagation. The complex quantity $\Gamma(k,\mathbf{n})$ depends on the details of the closing, i.e., on the approximations employed to solve the equations at depth $D+1$. Neglecting the couplings to DDOs at depth $D+2$ and using the Markovian and adiabatic approximations yields
@@
Following the practice common in models without explicit dissipation, in Eq. (35), we retain only the real parts of $\Gamma(k,\mathbf{n})$, i.e., we neglect the closing-induced renormalization of free-oscillation frequencies of DDOs at the terminal layer. We have checked that the imaginary parts of $\Gamma(k,\mathbf{n})$ in Eq. (35) very weakly affect the overall transport dynamics and, in particular, the long-time transport coefficient $\mu_{\mathrm{dc}}$. In undamped models, [64, 24] $\mathrm{Re}\>\Gamma_{\mathrm{MA}}(k,\mathbf{n})$ and $\mathrm{Re}\>\Gamma_{\mathrm{DR}}(k,\mathbf{n})$ are meaningful only in the infinite-chain limit $N\to\infty$, in which these can be evaluated analytically. For $\gamma_{0}\neq 0$, $\mathrm{Re}\>\Gamma_{\mathrm{MA}}(k,\mathbf{n})$ and $\mathrm{Re}\>\Gamma_{\mathrm{DR}}(k,\mathbf{n})$ can be computed for any finite $N$, on which they strongly depend. Although analytical expressions for $\mathrm{Re}\>\Gamma_{\mathrm{MA}}(k,\mathbf{n})$ and $\mathrm{Re}\>\Gamma_{\mathrm{DR}}(k,\mathbf{n})$ in the $N\to\infty$ limit are not available for $\gamma_{0}\neq 0$, we find that their long-chain limits are reached for $N$ much larger than typical chain lengths $N\sim 30$ we consider here. In practice, we find that computing Eqs. (37) and (38) on a chain containing $N_{\infty}=(10^{3}-10^{4})N$ sites provides us with the desired long-chain limits (although $N_{\infty}/N$ generally decreases with increasing $\gamma_{0}$).
@@
In Ref. 25, we have checked that both the Markovian-adiabatic [Eqs. (36) and (37)] and derivative-resum [Eq. (38)] schemes yield virtually identical transport dynamics for $\gamma_{0}=0$. Figures 2(b) and 2(c) confirm that the same holds in the overdamped regime $\gamma_{0}/\omega_{0}=1.7$. These also reveal that for sufficiently strong damping, the so-called time-nonlocal closing, [86] which sets $\Gamma(k,\mathbf{n})=0$ in Eq. (35), also yields stable long-time diffusive dynamics from which low-frequency dynamical mobility can be meaningfully extracted. This stands in sharp contrast to the undamped case, in which computations of long-time transport coefficients crucially require a closing term [25] because these cannot be accessed by extrapolating the results for small nonzero dampings to the zero-damping limit. [87] Although Figs. 2(b) and 2(c) suggest that the closing terms can be omitted if $\gamma_{0}$ is sufficiently large, all our results rely on the propagation of the closed DEOM. Using one and the same closing scheme throughout the physically reasonable range for $\gamma_{0}$ down to $\gamma_{0}=0$ enables a consistent and fair assessment of trends in transport dynamics with varying $\gamma_{0}$.
@@
Furthermore, Figs. 2(b) and 2(c) suggest that our DEOM closing enhances not only the convergence with respect to the maximum depth but also with respect to the chain length. In particular, the DEOM dynamics with $N=31$ and Markovian-adiabatic or derivative-resum closing are very similar to those employing a $\approx 50\%$-longer chain ($N=45$) and time-nonlocal closing. In the overdamped regime considered here, the diffusion-constant growth on intermediate timescales and the corresponding dynamical-mobility feature at $\omega\approx\omega_{0}$ reflect finite-size effects [see Figs. 2(b) and 2(c) for $N=31$]. Meanwhile, analogous features for weak damping are not artifacts due to insufficient chain length, [25] as discussed in greater detail in Sec. III.3.
@@
Finally, we prefer the Markovian-adiabatic [Eqs. (36) and (37)] to the derivative-resum [Eq. (38)] scheme because of its more favorable computational requirements. In particular, the Markovian-adiabatic scheme needs only $N$ quantities $\gamma_{k}$ computed on an $N_{\infty}$-site chain, while the derivative-resum scheme needs a much larger number of quantities $\Gamma_{\mathrm{DR}}(k,\mathbf{n})$, $N$ quantities for each DDO at depth $D$. Although all these quantities can be computed only once, before the r-DEOM propagation, our derivative-resum scheme requires much more computational time and computer memory to compute and store the closing terms than the Markovian-adiabatic scheme. It is for this reason that we employ the Markovian-adiabatic scheme in the following.
@@ III.2 Thermodynamic quantities
In addition, the carrier–bath interaction energy
@@
has a nonzero imaginary part, in contrast to the undamped case. [24, 25] Although Fig. 3(a) shows that both $\mathrm{Re}\>\langle H_{S}\rangle$ and $\mathrm{Re}\>\langle H_{S-B}\rangle$ grow quadratically with $\gamma_{0}$ (for sufficiently large $\gamma_{0}$), they are virtually equal to the corresponding values in the undamped case for physically plausible values of $\gamma_{0}$. Figure 3(b) shows that $\mathrm{Im}\>\langle H_{S}\rangle$ and $\mathrm{Im}\>\langle H_{S-B}\rangle$ grow linearly with $\gamma_{0}$ but remain quite small with respect to the corresponding real parts throughout the physically relevant range $\gamma_{0}/\omega_{0}\lesssim 2$. Therefore, the real parts of Eqs. (40) and (41) obtained using the i-DEOM remain representative of the kinetic and interaction energies, respectively.
@@ III.3 Finite-depth and finite-chain effects for weak damping
Figure 4: Dynamics of (a) the diffusion constant $\mathcal{D}$ and (b) the diffusion exponent $\alpha$ for different values of the chain length NN and maximum hierarchy depth $D$. (c) The dynamical-mobility profile $\mathrm{Re}\>\mu(\omega)$ for different values of $N$ and $D$. The insets in (a) and (b) zoom in the dynamics of $\mathcal{D}$ and $\alpha$, respectively, on intermediate-to-long timescales. The model parameters are $J=1,\omega_{0}=0.044,\lambda=0.336,T=0.175$, and $\gamma_{0}/\omega_{0}=0.05$. Note the logarithmic scale on the horizontal axes.
@@
Fixing $N=21$ or 31, we observe that transport properties computed for $D=4,5,6$ or $D=4,5$ are qualitatively similar, though there are some quantitative differences. In particular, the growth of $\mathcal{D}(t)$ beyond $\mathcal{D}(t_{\mathrm{min}})$ [Fig. 4(a)], where $t_{\mathrm{min}}\approx 1/\omega_{0}$, the overshoot of $\alpha(t)$ above unity on intermediate timescales [Fig. 4(b)], and the dynamical-mobility enhancement below $\omega_{0}$ [Fig. 4(c)] are more pronounced with increasing $D$. Meanwhile, increasing $D$ does not change either the position or shape of the displaced Drude peak [Fig. 4(c)], while quantitative changes in transport properties upon increasing $D$ are of the order of $10\%$[25] As DEOM computations for $D=5$ (let alone $D=6$) require significantly more computational resources than those employing $D=4$, we opt for $D=4$ as a good compromise between computational effort and accuracy. The relative accuracy with which the OSR [Eq. (42)] is satisfied is $\sim 10^{-4}$ for all depths considered.
@@
Fixing $D=4$ ($D=5$), we find that $\mathcal{D}(t),\alpha(t),$ and $\mathrm{Re}\>\mu(\omega)$ for $N=21$, 31, and 45 ($N=21$ and 31) are qualitatively and to a large extent quantitatively similar. Finite-chain effects on the dynamics of $\mathcal{D}$ and $\alpha$ come into play for $t\gtrsim\pi/\omega_{0}$ [see the insets of Figs. 4(a) and 4(b)]. We observe that increasing $N$ renders the increase in $\mathcal{D}(t)$ for $t\gtrsim t_{\mathrm{min}}$, the overshoot of $\alpha(t)$ above unity, and the dynamical-mobility enhancement below $\omega_{0}$ somewhat less pronounced. One might thus object to the relevance of these features in the long-chain limit. To refute such an objection, we note that the very same features have been observed in fully quantum [90] or quantum–classical [91, 21, 92, 93] approaches to carrier transport in models with local [90, 91] or nonlocal [21, 92, 93] interaction of a carrier with undamped [90, 91, 21, 92, 93] or moderately damped [21] vibrations. In particular, the quantum–classical study in Ref. 21 considered the same adiabaticity $\omega_{0}/J$ ratio as here and observed the same long-time (low-frequency) features in $\mathcal{D}(t)$ $\mathrm{Re}\>\mu(\omega)$] on chains with hundreds of sites. By assuming classical vibrational dynamics, which is appropriate for small $\omega_{0}/J$ and at the high temperature considered here, quantum–classical methods manage to reach what would be the $D\to\infty$ limit in the DEOM language. Keeping in mind that increasing $D$ beyond 4 renders the long-time (low-frequency) features more pronounced, their existence for small $\gamma_{0}$ can be considered as firmly established.
@@
The above observations strengthen the conclusions we reached in the dissipationless case, [25] firmly establishing the transient nature of the super-to-subdiffusive crossover and the low-frequency enhancement in the carrier’s optical response in the case of weakly damped vibrations moderately coupled to the carrier.